OpenStudy (bohotness):

Which value for x makes the sentence true? 748 + x = 1,269 A. 531 B. 2,117 C. 521 D. 2,017

OpenStudy (lizz123):

What does 748+531=

OpenStudy (lizz123):

@bohotness

OpenStudy (anonymous):

wot? a ss 91 cant do this lel

OpenStudy (bohotness):

1279

OpenStudy (lizz123):

Okay so now you know it can't be A.

OpenStudy (lizz123):

What does 748+521=

jigglypuff314 (jigglypuff314):

748 + x = 1,269 subtract 748 from both sides -748 -748 ----------------

OpenStudy (bohotness):

1269

OpenStudy (lizz123):

So finally you have the answer of 521

OpenStudy (lizz123):

C. is your answer

OpenStudy (bohotness):

THANKS

OpenStudy (lizz123):

your welcome